State whether the following are true or false. Justify your answer. 
The value of tan A is always less than 1

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

False

Justification​: In ΔMNC in which∠N = 90∘,
MN = 3, NC = 4 and MC = 5

Value of tan M = 4/3 which is greater than.
The triangle can be formed with sides equal to 3, 4 and hypotenuse = 5 as it will follow the Pythagoras theorem.
MC^2=MN^2+NC^2
5^2=3^2+4^2
25=9+16
25=25
Hence, The same can be observed.

Thanks